# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

### v2

@@ Abstract
We propose the X3Z3 Floquet code, a type of dynamical code with improved performance under biased noise compared to other Floquet codes. The enhanced performance is attributed to a simplified decoding problem resulting from a persistent symmetry under infinitely biased noise, which surprisingly exists in a code without constant stabilisers. Even if such a symmetry is allowed, we prove that a general dynamical code with two-qubit parity measurements cannot admit one-dimensional decoding graphs, a key feature resulting in the high performance of bias-tailored stabiliser codes. Despite this limitation, we demonstrate through our comprehensive numerical simulations that the symmetry of the X3Z3 Floquet code renders its performance under biased noise far better than several leading Floquet code candidates. Furthermore, to maintain high-performance implementation in hardwares without native two-qubit parity measurements, we introduce ancilla-assisted bias-preserving parity measurement circuits. Our work establishes the X3Z3 code as a prime quantum error-correcting code candidate, particularly for devices with reduced connectivity, such as the honeycomb and heavy-hexagonal architectures.
@@ I Introduction
Quantum error correction (QEC) [1, 2, 3, 4] should be understood as occurring both in space and time [5]. Taking advantage of the temporal dimension, Floquet codes [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], or more generally dynamical codes [21, 22, 23], form a large class of error-correcting codes, which can achieve competitive fault-tolerant performance [6, 8, 9, 10, 13, 12, 11, 14, 16, 15, 17] while reducing the weights of check measurements performed for the error correction. Several of these codes [7, 9, 18, 13] also benefit from being defined on a lattice with sparser connectivity than that for the surface code [24, 25]: each qubit is only connected to three other qubits. In architectures where two-qubit parity check measurements are native, Floquet codes could achieve higher thresholds [10, 8, 17] and lower space-time overheads [8, 10] than the surface code. Without requiring additional connectivities, this code can be deformed around defective components caused by highly noisy qubits or gates [15, 26]. Moreover, dynamical codes allow for implementations of arbitrary Clifford, and even some non-Clifford gates, through low-weight parity check measurements [21]. In addition, dynamical measurement schedules can result in certain errors being self-corrected [20].
@@
Although Floquet codes (in particular, the honeycomb code) have been studied under various noise models [9, 27], there have not been any Floquet codes that are specifically tailored for an improved performance under biased noise. A biased noise model is one in which a specific type of error, for example, phase errors, occurs more frequently than other errors, such as bit flip errors. This biased noise is typical to most quantum platforms, for example, superconducting fluxonium qubits [28], bosonic “cat” qubits [29, 30], spin-optical [31, 17], neutral atoms [32], quantum-dot spin qubits [33, 34] and Majorana qubits [35, 36].
@@
For enhanced performance, quantum-error-correcting codes need to be designed such that they possess symmetries that can be utilised to simplify the error-syndrome decoding problem given the noise structure [37, 38, 39, 40, 41, 42, 43, 44, 45]. While there have been a number of proposals for bias-tailored static codes [38, 39, 41, 45, 46, 40, 47, 41, 44, 42, 48, 49, 43, 50, 51], designing Floquet codes for high performance under biased noise is still an open problem. Owing to the experimental relevance of biased noise and given the ease of implementation of Floquet codes which require only two-qubit parity measurements, it is therefore imperative to tailor Floquet codes for biased noise and study how the performance of such dynamical codes can be improved.
@@
In this paper, we present the X3Z3 bias-tailored Floquet code, a Clifford-deformed [45] version of the Calderbank–Shor–Steane (CSS) Floquet code [18, 13]. Despite not having a fixed stabiliser group (as static codes have), the X3Z3 Floquet code surprisingly still possesses a symmetry under infinitely biased noise that allows for a simplified decoding in biased noise regimes. We perform an in-depth study of this code, along with the CSS Floquet code [18, 13], and two types of honeycomb codes: one introduced by Hastings and Haah [6] and the other by Gidney et al. [8]. We simulate all codes under biased-noise models, and find that the X3Z3 code has the best performance. Using a matching decoder, we find that, as the noise changes from fully depolarising to pure dephasing, the X3Z3 Floquet code threshold increases from $1.13\%$ to $3.09\%$ under a code-capacity noise model and increases from $0.76\%$ to $1.08\%$ under a circuit-level noise model mimicking hardwares with noisy direct entangling measurements. Furthermore, we show that its sub-threshold performance is also substantially better under biased noise than other Floquet codes.
@@
Compared to its static counterparts, the X3Z3 Floquet code has an advantage that it can be realised using only two-qubit parity check measurements. This makes it particularly suitable for devices with constrained connectivities, such as the honeycomb and heavy-hexagonal lattice (currently IBM’s preferred superconducting-qubit architecture) [52, 53]. Moreover, we demonstrate that the two-qubit parity measurements of the Floquet code can be performed in a bias-preserving way even in hardware without direct entangling measurements, thus enabling high performance implementation in such devices.
@@
To explain the performance of our bias-tailored Floquet code, we argue that other dynamical codes defined on the same architecture, and built from two-body measurements, would likely not have drastically improved performance compared to the X3Z3 Floquet code. To support this argument, we prove that decoding graphs of such dynamical codes under infinitely biased noise have connectivities that are too high for the decoding problem to be reduced to a simple decoding of repetition codes, as is the case for static codes [41, 45]. This can be understood as resulting from the fact that error syndromes of dynamical codes possess less symmetry than their static code counterparts.
@@
The paper is laid out as follows. In Sec. II, we review the basics of Floquet codes together with two widely studied examples: honeycomb and CSS Floquet codes. Readers who are already familiar with Floquet codes can skip directly to Sec. III where we discuss our X3Z3 Floquet code which is obtained by bias-tailoring the CSS Floquet code. Crucially, we show that there exists a persistent symmetry in the code’s error syndrome under pure dephasing noise that allows for a simplified decoding. Subsequently, we introduce ancilla-assisted bias-preserving parity measurement circuits that allow for high-performance code implementation in devices without native entangling measurements. In Sec. IV, we present our simulation methods and noise models. In Sec. V, we then provide our simulation results for all of the four Floquet codes studied. To explain the code performance, in Sec. VI we prove that a class of dynamical codes on the honeycomb lattice has high-connectivity decoding graphs. We also prove that dynamical codes with only two-qubit parity check measurements cannot have repetition-code-like decoding graphs, as is the case with static codes. Finally, we conclude and present future research directions in Sec. VII. In the Appendices, we give a more detailed review of the basics of honeycomb and CSS Floquet codes, and present the details of our noise models, numerical simulations, and theorems.
II Background
@@ II.1 Dynamical and Floquet codes
Figure 1: (a) CSS and (b,c) X3Z3 Floquet codes. (a,b) Left: Hexagonal lattice with qubits sitting on vertices and opposite boundaries identified. Plaquettes and edges are assigned one of three colours [red, green, blue; examples are shown in (a)] and one of two types (X- or Z-type for CSS and A- or B-type for X3Z3). (a) Right: CSS Floquet code measurement schedule. (b) Left: The X3Z3 Floquet code is obtained from the CSS code via Hadamard gates applied to shaded domains’ qubits. Right: Plaquette and edge (check) operators are A- or B-type, depending on their support basis on the shaded/unshaded domains. (c) The X3Z3 Floquet code measurement schedule. Arrows indicate the type and colour of the edge operator measured at each step, where the edges just measured (members of the ISG) are highlighted in the lattice. Uncoloured plaquettes host only a single type of stabiliser, either A or B, indicated by the letters in the plaquettes, while coloured plaquettes host both A- and B-type stabilisers in the ISG. One set of anti-commuting logical operators is shown by yellow and brown strings, where their $X$ ($Z$) bases on the qubit supports are highlighted in red (blue). The other set (not shown) is similar to the set shown but offset by three measurement subrounds and with an $X\leftrightarrow Z$ interchange of the qubit supports’ bases.
@@
We define qubits as residing on the vertices of the lattice and error-correction will be performed by measuring two-qubit operators defined on edges of the lattice (i.e., acting on qubits incident to a given edge). Each edge is coloured the same as the plaquettes it connects [see e.g., the coloured edges in the middle right of the lattice in Fig. 1(a)]. We perform error-correction by measuring edge operators in a particular sequence. After any given subround of edge measurements, the system will be in the $+1$-eigenstate of the operators in an “instantaneous stabiliser group” (ISG), which will change at each time step. The ISG at time step $t+1$ is defined as
$\begin{split}\text{ISG}_{t+1}=\langle S,\pm M\;|\;M\in\mathcal{M}_{t+1},\;S\in\text{ISG}_{t},\\ \text{such that }[S,M]=0,\;\forall M\in\mathcal{M}_{t+1}\rangle.\end{split}$ (1)
@@
The logical operators at time $t$ are given by $\mathcal{C}(\mathrm{ISG}_{t})/\text{ISG}_{t}$, where $\mathcal{C}(\text{ISG}_{t})$ is the centraliser of ISGt, i.e., the group of Pauli operators commuting with all $S\in\text{ISG}_{t}$. A (potentially trivial) logical operator “representative” is some member of $\mathcal{C}(\text{ISG}_{t})$. Each nontrivial logical operator representative for the codes considered will be a string-like Pauli operator at each time step [see Fig. 1(c)]. To avoid anti-commuting with the next-subround edge measurements, certain check measurement results along a logical operator’s path have to multiplied into that logical operator. Hence, the logical operators will evolve from one time step to the next.
@@
Having discussed the general idea of Floquet codes, we now briefly review the commonly studied examples of the code: two variants of the honeycomb code [6, 8], whose performance will also be investigated under biased noise, and the CSS Floquet code [18, 13]. We will later modify the CSS Floquet code to achieve the bias-tailored X3Z3 Floquet code. More details of the honeycomb and CSS Floquet codes are presented in Appendix A.
@@ II.2 Honeycomb codes
We begin by first discussing the honeycomb codes. The first variant is due to Gidney et al. [8], which we call the P6 Floquet code, since its plaquette operators are six-body operators of the form $P^{\otimes 6}$ for $P=X,Y,Z$. We define edge operators of three types: on red edges we define an $XX$ operator, on green edges a $YY$ operator and on blue edges a $ZZ$ operator. We measure edge operators in the periodic sequence $r\rightarrow g\rightarrow b$. When this code is defined with periodic boundary conditions it stores two logical qubits (it is equivalent to the toric code concatenated with a two-qubit repetition code at each time step [6]). The code’s logical operators evolve through the measurement cycle (see Appendix A.1). While the measurement sequence has period 3, the logical operators only return to their initial values (up to signs) with period 6.
We define one stabiliser operator for each plaquette, where red, green, and blue plaquettes host the $X^{\otimes 6}$, $Y^{\otimes 6}$, and $Z^{\otimes 6}$ operators, respectively. These plaquette operators are inferred from edge measurements in two consecutive subrounds. Detectors are formed from consecutive plaquette operator measurements.
The second honeycomb code variant, which we call the XYZ2 honeycomb code, is due to Hastings and Haah [6]. It differs from the P6 code by single-qubit Clifford rotations acting on the qubits. While the XYZ2 code’s edges are still coloured the same as the P6 edges, the XYZ2 code’s edge operators have their Pauli bases defined according to their orientations within each T junction, i.e., horizontal edges are of $Z$ type while the $X$ and $Y$ checks are respectively those edges which are 90 clockwise and counter-clockwise rotated from the horizontal edges (see Appendix A.1). All plaquettes have the same stabiliser operator, i.e., the XYZ2 operator. While the logical operators of the XYZ2 code have the same qubit supports as those of the P6 code, the qubit support bases of the XYZ2 logical operators are not uniform throughout, but involve $X$, $Y$, and $Z$ Paulis [6]. We note that even though these honeycomb code variants have been studied by several works in the literature (e,g., Refs. [6, 8, 7, 9]), there have not been any studies comparing the performance of these two codes under biased noise.

### v3

@@ Abstract
We propose the X3Z3 Floquet code, a dynamical code with improved performance under biased noise compared to other Floquet codes. The enhanced performance is attributed to a simplified decoding problem resulting from a persistent stabiliser-product symmetry, which surprisingly exists in a code without constant stabilisers. Even if such a symmetry is allowed, we prove that general dynamical codes with two-qubit parity measurements cannot admit one-dimensional decoding graphs, a key feature responsible for the high performance of bias-tailored stabiliser codes. Despite this, our comprehensive simulations show that the symmetry of the X3Z3 Floquet code renders its performance under biased noise far better than several leading Floquet codes. To maintain high-performance implementation in hardware without native two-qubit parity measurements, we introduce ancilla-assisted bias-preserving parity measurement circuits. Our work establishes the X3Z3 code as a prime quantum error-correcting code, particularly for devices with reduced connectivity, such as the honeycomb and heavy-hexagonal architectures.
@@ I Introduction
Quantum error correction (QEC) [1, 2, 3, 4] should be understood as occurring both in space and time [5]. Taking advantage of the temporal dimension, Floquet codes [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], or more generally dynamical codes [21, 22, 23], form a large class of quantum error-correcting codes, which can achieve competitive fault-tolerant performance [6, 8, 9, 10, 13, 12, 11, 14, 16, 15, 17] while reducing the weights of check measurements performed for the error correction. Several of these codes [7, 9, 18, 13] also benefit from being defined on a lattice with sparser connectivity than that for the surface code [24, 25]: each qubit is only connected to three other qubits. In architectures where two-qubit parity check measurements are native, Floquet codes could achieve higher thresholds [10, 8, 17] and lower space-time overheads [8, 10] than the surface code. Without requiring additional connectivity, such codes can be deformed around defective components caused by highly noisy qubits or gates [15, 26]. Moreover, dynamical codes allow for implementations of arbitrary Clifford, and even some non-Clifford gates, through low-weight parity check measurements [21]. In addition, dynamical measurement schedules can result in certain errors being self-corrected in the Floquet-Bacon-Shor code [20] – this code was recently demonstrated in superconducting qubit experiments [27].
@@
Although Floquet codes (in particular, the honeycomb code) have been studied under various noise models [9, 28], there have not been any Floquet codes that are specifically tailored for an improved performance under biased noise. A biased noise model is one in which a specific type of error, for example, phase errors, occurs more frequently than other errors, such as bit flip errors. This biased noise is typical to most quantum platforms, for example, bosonic “cat” qubits [29, 30], spin-optical [31, 17], neutral atoms [32], quantum-dot spin qubits [33, 34] and Majorana qubits [35, 36]. In these platforms, biased noise can arise due to different predominant error mechanisms. For example, by increasing the photon number of the resonators, the cat qubits [29, 30] can be made exponentially protected against the bit-flip error at the expense of only a linear increase in the phase-flip error. For the spin qubits in both spin-optical [17] and quantum dot architectures [33, 34], the noise is predominantly dephasing due to the short spin coherence time, $T_{2}$. Moreover, the two-qubit gates of quantum-dot spin qubits have also been experimentally shown to exhibit phase-biased noise which is caused by the $T_{2}$ dephasing, non-Markovian error sources, AC-Stark shift, and calibration errors [34]. In neutral-atom qubits, detailed modeling of the experiment shows that the two-qubit gate noise is predominantly phase-biased due to the short Rydberg state $T_{2}^{*}$ lifetime caused by finite atomic temperature and light-shift fluctuations of the lasers [32]. Lastly, for Majorana qubits, the noise is expected to be biased due to residual coupling between the Majoranas and thermal fluctuations that excite the qubit to an above-gap quasiparticle state [37].
@@
For enhanced performance, quantum-error-correcting codes need to be designed such that they possess symmetries that can be utilised to simplify the error-syndrome decoding problem given the noise structure [38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. While there have been a number of proposals for bias-tailored static codes [48, 39, 40, 42, 46, 49, 41, 50, 42, 45, 43, 51, 52, 44, 53, 54, 55, 47], designing Floquet codes for high performance under biased noise is still an open problem. Owing to the experimental relevance of biased noise and given the ease of implementation of Floquet codes, which require only two-qubit parity measurements that can be implemented even in architectures with sparse connectivity, it is therefore imperative to tailor Floquet codes for biased noise and study how the performance of such dynamical codes can be improved.
@@
In this paper, we present the X3Z3 bias-tailored Floquet code, a Clifford-deformed [46] version of the Calderbank–Shor–Steane (CSS) Floquet code [18, 13]. Despite not having a fixed stabiliser group (as static codes have), remarkably the X3Z3 Floquet code possesses a stabiliser-product symmetry under infinitely phase-biased noise, simplifying decoding in biased noise regimes. We perform an in-depth study of this code, along with the CSS Floquet code [18, 13], and two types of honeycomb codes: one introduced by Hastings and Haah [6] and the other by Gidney et al. [8]. We simulate all codes under biased-noise models, and find that the X3Z3 code has the best performance. Using a matching decoder, we find that, as the noise changes from fully depolarising to pure dephasing, the X3Z3 Floquet code threshold increases from $1.13\%$ to $3.09\%$ under a code-capacity noise model and increases from $0.76\%$ to $1.08\%$ under a circuit-level noise model mimicking hardware with noisy direct entangling measurements. Furthermore, we show that its sub-threshold performance is also substantially better under biased noise than other Floquet codes.
@@
Compared to its static counterparts, the X3Z3 Floquet code has an advantage that it can be realised using only two-qubit parity check measurements. This makes it particularly suitable for devices with constrained connectivity, such as the honeycomb and heavy-hexagonal lattice (currently IBM’s preferred superconducting-qubit architecture) [56, 57]. Moreover, we demonstrate that the two-qubit parity measurements of the Floquet code can be performed in a bias-preserving way even in hardware without direct entangling measurements, thus enabling high performance implementation in such devices. Crucially, we show that even in the presence of mid-gate errors which degrade noise bias in the target qubits of conventional CNOT gates [58], two of the three required parity-check circuits, built using these conventional gates, can be made fully $Z$-bias preserving, i.e., sustain the $Z$ noise bias on both data qubits, while the other one is partially phase-bias preserving, i.e., protects the $Z$ noise bias in only one of the data qubits.
@@
We argue that other dynamical codes defined on the same architecture, and built from two-body measurements, would likely not have drastically improved performance compared to the X3Z3 Floquet code. To support this argument, we prove that decoding graphs of such dynamical codes under infinitely phase-biased noise have connectivities that are too high for the decoding problem to be reduced to a simple decoding of repetition codes, as is the case for static codes [42, 46]. This can be understood as resulting from the fact that error syndromes of dynamical codes possess less symmetry than their static code counterparts.
@@
The paper is laid out as follows. In the Secs. II.1II.3, we first review the basics of Floquet codes together with two widely studied examples: honeycomb and CSS Floquet codes. Readers who are already familiar with Floquet codes can skip directly to Sec. II.4 in which we discuss our X3Z3 Floquet code. In Sec. II.5, we show that there exists a persistent symmetry in the code’s error syndrome under pure dephasing noise that allows for a simplified decoding. Subsequently, in Sec. II.6 we introduce ancilla-assisted bias-preserving parity measurement circuits that allow for high-performance code implementation in devices without native entangling measurements. In Secs. II.7 and II.8, we provide the simulation results for memory experiments. In Sec. II.9, we give two theorems showing that the decoding hypergraphs of dynamical codes cannot be reduced to 1D graphs. Finally, we conclude and present future research directions in Sec. III. In Sec. IV, we provide the description of our noise models, details of numerical simulations and proofs of the no-go theorems. In the Supplementary Information, we provide a more thorough review of the basics of honeycomb and CSS Floquet codes, details of our parity check circuits, plots for determining thresholds, details of hyperedges in the honeycomb codes, and results for Floquet codes with elongated dimension and twisted periodic boundary conditions.
II Results
@@ II.1 Dynamical and Floquet codes
Figure 1: X3Z3 and CSS Floquet codes on the hexagonal lattice with their stabilisers, checks, and logical operators. (a) CSS and (b,c) X3Z3 Floquet codes. (a,b) Left: Hexagonal lattice with qubits sitting on vertices and opposite boundaries identified. Plaquettes and edges are assigned one of three colours [red, green, blue; examples are shown in (a)] and one of two types (X- or Z-type for CSS and A- or B-type for X3Z3). (a) Right: CSS Floquet code measurement schedule. (b) Left: The X3Z3 Floquet code is obtained from the CSS code via Hadamard gates applied to shaded domains’ qubits. Right: Plaquette and edge (check) operators are A- or B-type, depending on their support basis on the shaded/unshaded domains. (c) The X3Z3 Floquet code measurement schedule. Arrows indicate the type and colour of the edge operator measured at each step, where the edges just measured (members of the ISG) are highlighted in the lattice. Uncoloured plaquettes host only a single type of stabiliser, either A or B, indicated by the letters in the plaquettes, while coloured plaquettes host both A- and B-type stabilisers in the ISG. One set of anti-commuting logical operators is shown by yellow and brown strings, where their qubit supports are depicted using big circles and stars, respectively, with the $X$ ($Z$) bases highlighted by red (blue) colouring. The other set (not shown) is similar to the set shown but offset by three measurement subrounds and with an $X\leftrightarrow Z$ interchange of the qubit supports’ bases.
We define qubits as residing on the vertices of the lattice and error-correction is performed by measuring two-qubit operators defined on edges of the lattice (i.e., acting on qubits incident to a given edge). Each edge is coloured the same as the plaquettes it connects [see e.g., the coloured edges in the middle right of the lattice in Fig. 1(a)]. We perform error-correction by measuring edge operators in a particular sequence. After any given subround of edge measurements, the system will be in the $+1$-eigenstate of the operators in an “instantaneous stabiliser group” (ISG), which will change at each time step. The ISG at time step $t+1$ is defined as
$\text{ISG}_{t+1}=$ $\,\langle S,\pm M\;|\;M\in\mathcal{M}_{t+1},\;S\in\text{ISG}_{t},$
$\quad\text{such that }[S,M]=0,\;\forall M\in\mathcal{M}_{t+1}\rangle.$ (1)
@@
The logical operators at time $t$ are given by $\mathcal{C}(\mathrm{ISG}_{t})/\text{ISG}_{t}$, where $\mathcal{C}(\text{ISG}_{t})$ is the centraliser of ISGt, i.e., the group of Pauli operators commuting with all $S\in\text{ISG}_{t}$. A (potentially trivial) logical operator “representative” is some member of $\mathcal{C}(\text{ISG}_{t})$. Each lowest-weight nontrivial logical operator representative for the codes considered will be a string-like Pauli operator at each time step [see Fig. 1(c)]. To avoid anti-commuting with the next-subround edge measurements, certain check measurement results along a logical operator’s path have to be multiplied into that logical operator. Hence, the logical operators will evolve from one time step to the next. As an illustration, take for example the vertical logical operator immediately after the $g$B checks are applied, i.e., the operator on the yellow string in the top-centre panel of Fig. 1(c). This operator is obtained by multiplying the vertical green-B checks on the vertical logical operator’s path [the two green edges on the yellow path shown in the top-centre panel of Fig. 1(c)] into the previous time-step vertical logical operator [the operator shown by the yellow string in the top-left panel of Fig. 1(c)]. Note that after the multiplication, the logical operator commutes with the next check measurements, which are the blue-A edge measurements.
@@
Having discussed the general idea of Floquet codes, we now briefly review the commonly studied examples: two variants of the honeycomb code [6, 8] and the CSS Floquet code [18, 13]. We will later modify the CSS Floquet code to achieve the bias-tailored X3Z3 Floquet code. More details of the honeycomb and CSS Floquet codes are presented in Supplementary Sec. I of the Supplementary Information [61].
@@ II.2 Honeycomb codes
We begin by first discussing the honeycomb codes. The first variant we review is due to Gidney et al. [8], which we call the P6 Floquet code, since its plaquette operators are six-body operators of the form $P^{\otimes 6}$ for $P=X,Y,Z$. We define edge operators of three types: on red edges we define an $XX$ operator, on green edges a $YY$ operator and on blue edges a $ZZ$ operator. We measure edge operators in the periodic sequence $r\rightarrow g\rightarrow b$. When this code is defined with periodic boundary conditions it stores two logical qubits (it is equivalent to the toric code concatenated with a two-qubit repetition code at each time step [6]). The code’s logical operators evolve through the measurement cycle (see Supplementary Sec. IA of the Supplementary Information [61]). While the measurement sequence has period 3, the logical operators only return to their initial values (up to signs) with period 6.
We define one stabiliser operator for each plaquette, where red, green, and blue plaquettes host the $X^{\otimes 6}$, $Y^{\otimes 6}$, and $Z^{\otimes 6}$ operators, respectively. These plaquette operator eigenvalues are inferred from edge measurements in two consecutive subrounds. Detectors are formed from consecutive plaquette operator measurements.
The second honeycomb code variant, which we call the XYZ2 honeycomb code, is due to Hastings and Haah [6]. It differs from the P6 code by single-qubit Clifford rotations acting on the qubits. While the XYZ2 code’s edges are still coloured the same as the P6 edges, the XYZ2 code’s edge operators have their Pauli bases defined according to their orientations within each T junction, i.e., horizontal edges are of $Z$ type while the $X$ and $Y$ checks are respectively those edges which are 90 clockwise and counter-clockwise rotated from the horizontal edges (see Supplementary Sec. IA of the Supplementary Information [61]). All plaquettes have the same stabiliser operator, i.e., the XYZ2 operator. While the logical operators of the XYZ2 code have the same qubit supports as those of the P6 code, the qubit support bases of the XYZ2 logical operators are not uniform throughout, but involve $X$, $Y$, and $Z$ Paulis [6]. We note that even though these honeycomb code variants have been studied by several works in the literature (e,g., Refs. [6, 8, 7, 9]), there have not been any studies comparing the performance of these two codes under biased noise.